OpenStudy (anonymous):

Geometry! If the sun of the interior angles of a polygon is 540 degrees then how many sides does the polygon have?

OpenStudy (anonymous):

I figured it out...The sum of the interior angles of a polygon is equal to 2 less than the number of sides the polygon has multiplied by 180. This can be expressed in an equation as: a=180(s-2) where a equals the sum of interior angles and s equals the number of sides the polygon has. If we plug 540 in as the 'a' value, we find that: Ads by Video PlayerAd Options 540=180(s-2) We can then divide both sides of the equation by 180: 3=s-2 Finally, we can add 2 to both sides of the problem: 5=s Therefore we find that the polygon has 5 sides.